OpenStudy (openstudyuser):

Painting: Need help with getting equations! Michael Thomas, the manager of a paint store, is mixing paint for a spring sale. There are 32 units of yellow dye, 54 units of brown dye, and an unlimited supply of base paint available. Mr.Thomas plans to mix as many gallons as possible of Autumn Wheat and Harvest Brown paint. Each gallon of Autumn Wheat requires 4 units of yellow dye, and 1 unit of brown dye. Each gallon of Harvest Brown paint requires 1 unit of yellow dye and 6 units of brown dye. Find the maximum number of gallons of paint that Mr.Thomas can mix.

OpenStudy (openstudyuser):

Yellow Dye= x Brown dye=y

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

So first of all \(x\leq 32 \) and \(y \leq 54\)

OpenStudy (openstudyuser):

So when he says there are unlimited, we put greater than or equal to?

OpenStudy (anonymous):

Well the dye is mixed with the base paint. The point is we won't run out of base paint, but will run out of dye.

OpenStudy (anonymous):

We can ignore the base paint, basically.

OpenStudy (openstudyuser):

So how did you get the greater than or equal to sign for the both of them?

OpenStudy (anonymous):

\(A = 4x+y\) and \(H=1x+6y\)

OpenStudy (anonymous):

Well, it's okay NOT to run out of dye.

OpenStudy (anonymous):

So if we use less than how much we have, we are still ok.

OpenStudy (openstudyuser):

For this one, we have two major equations?

OpenStudy (anonymous):

yeah but we may combine them too.

OpenStudy (openstudyuser):

Isnt Autumn Wheat the yellow paint and Harvest brown the brown one right?

OpenStudy (anonymous):

Yeah.

OpenStudy (openstudyuser):

"Brown paint. Each gallon of Autumn Wheat requires 4 units of yellow dye, and 1 unit of brown dye. Each gallon of Harvest Brown paint requires 1 unit of yellow dye and 6 units of brown dye" Thats the stuff i was mainly confused on.

OpenStudy (anonymous):

That is represented with my A, and H equations.

OpenStudy (openstudyuser):

Basically, what they are saying is that to make a Autumn Wheat paint, they need to mix 4 of the yellow and 1 of the brown. And to make Harvest Brown, they need 1 yellow and 6 brown?

OpenStudy (anonymous):

Yes.

OpenStudy (openstudyuser):

You would have to graph those A and H equations?

OpenStudy (anonymous):

I don't think so.

OpenStudy (anonymous):

We have many unknowns.

OpenStudy (openstudyuser):

Yea, cuz we need one more equation to graph, if we only graph, \[x \le32\] and \[y \le54\] Wouldnt that make the graph unbounded?

OpenStudy (anonymous):

We don't know how many of A or H we'll make, so we need a quantity for that. The total gallons, \(g\) is gonna be: \( g= a(4x+y) + h(x+6y)\)

OpenStudy (anonymous):

Well we also know we can't get into negatives, so it's more like \[0\leq x \leq 32\]\[0\leq y \leq 54 \]

OpenStudy (anonymous):

But that's a lot of potential points to consider, soooo...

OpenStudy (openstudyuser):

Iam lost. lol.

OpenStudy (anonymous):

I think it's one of those problems where you just have to try a bunch of points out.

OpenStudy (openstudyuser):

In which equations?

OpenStudy (anonymous):

Heh, that's the thing, even when you know what point you're going to use, you still have to decide whether you want to make Autumn or Harvest... let me think on this a bit.

OpenStudy (openstudyuser):

Okay, im just very confused on this concept lol.

OpenStudy (anonymous):

I mean, if we only make Autumn paint, how many gallons can we make, if we only make Harvest paint, how many can we make? If we try for half and half, how many do we make?

OpenStudy (openstudyuser):

A whole? lol i dunno?

OpenStudy (anonymous):

Hmmm, actually I think I might has confused myself too, and messed up one of the equations.

OpenStudy (anonymous):

Okay, so if we only made autumn we'd make 8, with a ton of brown dye remaining. If we only make harvest we have 9 with a ton of yellow dye remaining.

OpenStudy (anonymous):

Catch my drift?

OpenStudy (openstudyuser):

Ummmmm...no not really lol.

OpenStudy (anonymous):

AW pain takes 4 yellow, we only have 32 yellow so 32/4 = 8 is the most we could make of it.

OpenStudy (anonymous):

Doing so would only take up 8 * 1 = 8 brown, and we'd have 54 - 8 = 46 brown left.

OpenStudy (anonymous):

For HB, it takes 6 brown. 54 / 6 = 9 we could make total, with 32 - 9 = 23 yellow pain remaining.

OpenStudy (openstudyuser):

If that were the case wouldnt\[x \ge32\] and \[y \ge54\]?

OpenStudy (anonymous):

No, they are less then or equal.

OpenStudy (anonymous):

saying they are greater than or equal would mean we could use as much as we wanted.

OpenStudy (openstudyuser):

But wouldnt we have limited paints?

OpenStudy (anonymous):

we have limited dyes, but unlimited base paint. I'm talking about dye.

OpenStudy (openstudyuser):

So shouldnt it be less than equal to?

OpenStudy (anonymous):

it is

OpenStudy (anonymous):

Hold on, I figured out what we've been doing wrong.... Our variables should have been how much AW and HB we make.

OpenStudy (openstudyuser):

So what would it have to equal to?

OpenStudy (dumbcow):

i know this is a little late but here is a nice concise solution: Let A be gallons of Autumn yellow , H be gallons of Harvest brown -you need 2 restraint equations: 1 for yellow dye and 1 for brown dye yellow: \[4A + H \le 32\] where "4A+H" represents total amount of yellow dye used brown: \[A + 6H \le 54\] -Graph these 2 equations to see possible region  total gallons = A+H solve for point where lines intersect to obtain maximum A=6 H=8 --> max gallons = 14